OpenStudy (anonymous):

Approximately 15% of Americans have a blood type of B. The local blood bank has 80 donors stop by on a given day (which can be considered a random sample from the population). What is the probability that between 15% and 20% of the donors on a given day have a blood type of B? Be sure to state and justify your assumptions.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

This is one of the other ones I need help with

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

I think it's a binomial distribution problem, but I'm double checking to make sure

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

ok I've confirmed it's a binomial distribution problem, here is the calculator we'll use http://stattrek.com/online-calculator/binomial.aspx

jimthompson5910 (jim_thompson5910):

the first two boxes will have these two values 0.15 80

OpenStudy (anonymous):

Ok1

OpenStudy (anonymous):

I did

jimthompson5910 (jim_thompson5910):

the next box will have the value of 16 why 16? because 20% of 80 is 16 this is the right endpoint of the interval

jimthompson5910 (jim_thompson5910):

so you'll have 0.15 80 16 hit enter

OpenStudy (anonymous):

kk i did

jimthompson5910 (jim_thompson5910):

look for: Cumulative Probability: P(X <= 16) make sure there's a line under the <

OpenStudy (anonymous):

it says binomial Probability: P(X = 16) : 0.053822345775557

jimthompson5910 (jim_thompson5910):

look for the value next to: Cumulative Probability: P(X <= 16)

OpenStudy (anonymous):

0.916288694306713

jimthompson5910 (jim_thompson5910):

good, keep this number in mind

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

now take 15% of 80 to get 0.15*80 = 12 then move to the next integer below that to get 11 now replace the 16 with 11, so you should have 0.15 80 11 hit enter and read off the number next to "Cumulative Probability: P(X <= 11) "

OpenStudy (anonymous):

.4522??

OpenStudy (anonymous):

0.452190494037699?

jimthompson5910 (jim_thompson5910):

good, you rounded, but that's ok

jimthompson5910 (jim_thompson5910):

now subtract this from the previous value you got 0.916288694306713-0.452190494037699 = 0.46409820026901

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that the answer?

jimthompson5910 (jim_thompson5910):

this is the probability that between 12 and 16 people (both inclusive) will have blood type B

OpenStudy (anonymous):

0.916288694306713-0.452190494037699 = 0.46409820026901

jimthompson5910 (jim_thompson5910):

12 is 15% of 80 16 is 20% of 80 so the probability that between 15% and 20% of the donors on a given day have a blood type of B is roughly 0.46409820026901 and you can round that however you need to

OpenStudy (anonymous):

so the answer: so the probability that between 15% and 20% of the donors on a given day have a blood type of B is roughly 0.46409820026901

jimthompson5910 (jim_thompson5910):

yeah you'll have to show how you got it of course

OpenStudy (anonymous):

i only have like one more essay then like multple choice ones, it will only take a litter. ill post it now ! thanks!

jimthompson5910 (jim_thompson5910):

basically using this idea P(k <= X <= m) = P(X <= m) - P(X<= k-1) for binomial distributions

OpenStudy (anonymous):

kk posting my new one